# On endomorphism algebras of silting complexes over hereditary abelian categories

### v2

@@ 3.1. Idempotent quotients
For a triangulated category $\mathcal{T}$ and a subcategory $\mathcal{X}$, we denote by $\mathcal{X}]$ the ideal of $\mathcal{T}$ consisting of the morphisms in $\mathcal{T}$ which factorize through objects in $\mathcal{X}$.
@@
Note that $\langle e_{D}\rangle$ consists of those morphisms in $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ that factor through objects in $\operatorname{\mathsf{add}}\nolimits D$. In other words, $\langle e_{D}\rangle=[\operatorname{\mathsf{add}}\nolimits D](T,T)$. Let $f\in[\operatorname{\mathsf{thick}}D](T,T)$, and assume $f$ factors as $T\xrightarrow{a}X\xrightarrow{b}T$ with $X\in\operatorname{\mathsf{thick}}D$. By Lemma 2.4, we have
@@
Let $A$ be an endomorphism algebra of a silting complex over a hereditary abelian category, and $e$ be an idempotent element of $A$.
@@
    Then $A/\langle e\rangle$ is also an endomorphism algebra of a silting complex over a hereditary abelian category.
@@
    Moreover, if $A$ is a quasi-silted algebra, then $A/\langle e\rangle$ is also a quasi-silted algebra.
@@
Let $\mathcal{H}$ be a hereditary abelian category, and $T\in{\mathcal{D}}^{b}(\mathcal{H})$ a basic silting object such that $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$. Let $T=D\oplus N$ and $e_{D}$ be the idempotent of $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ corresponding to $D$. It suffices to show that $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)/\langle e_{D}\rangle$ is isomorphic to an endomorphism algebra of a basic silting complex over a hereditary abelian category.
@@
where $f$ is a minimal right $\mathcal{S}$-approximation of $N$. It follows that $S_{N}\in\mathcal{S}^{\perp}$ by Lemma 2.2. On the other hand, by Lemma 2.4, it is not hard to see that $S\in\mathcal{S}_{D}^{\leq 0}:=\bigcup_{l\geq 0}\operatorname{\mathsf{add}}\nolimits D\ast\cdots\ast\operatorname{\mathsf{add}}\nolimits D[l]$.
@@
Now assume $T$ is 2-term (i.e. , $T$ satisfies (S1)) in ${\mathcal{D}}^{b}(\mathcal{H})$. By Proposition 2.8, we may identify ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ with $\mathcal{S}^{\perp}$ via the embedding $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\hookrightarrow{\mathcal{D}}^{b}(\mathcal{H})$. It suffices to verify that $S_{N}$ is 2-term in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$.
@@
Let $A$ be a finite dimensional algebra over an algebraically closed field, and $e$ be an idempotent element in $A$. By Theorem 3.1 and Theorem 3.3 (b), if $A$ is shod, then $A/\langle e\rangle$ is also shod.
@@ 3.2. Idempotent subalgebras
Let $A$ be an endomorphism algebra of a basic silting complex over a hereditary category, and $e$ be an idempotent element in $A$.
@@
    The subalgebra $eAe$ is also an endomorphism algebra of a basic silting complex over a hereditary category.
@@
    If $A$ is a quasi-silted algebra, then $eAe$ is also a quasi-silted algebra.
@@
    If $A$ is a quasi-tilted algebra, then $eAe$ is also a quasi-tilted algebra.
@@
(a) Assume $A=\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$, where $\mathcal{H}$ is a hereditary abelian category, and $T$ is a basic silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$. Assume $T=U\oplus N$ and $e=e_{N}$ is the idempotent corresponding to $N$. We first note that $e_{N}(\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T))e_{N}\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N)$. By Theorem 2.17, there exists a presilting complex $D$, such that $D\oplus N$ is a basic silting object and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$. In addition, $N$ is a silting object in $(\operatorname{\mathsf{thick}}D)^{\perp}$. By Proposition 2.7, $D^{\perp_{\mathbb{Z}}}$ is a hereditary abelian category. By Proposition 2.8, we identify $(\operatorname{\mathsf{thick}}D)^{\perp}$ with ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. Then $N$ is a basic silting complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$, and $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N)\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{D^{\perp_{\mathbb{Z}}}})}(N)$.
@@
Let $A$ be a finite dimensional algebra over an algebraically closed field, and $e$ be an idempotent element in $A$. By Theorem 3.1 and Theorem 3.5 (b), if $A$ is shod, then $eAe$ is also shod.
@@
The statements in parts (b) and (c) of Theorem 3.5 was established in [AC]. Our method is different with the one in [AC].
@@
Let $Z\in\operatorname{\mathsf{mod}}\nolimits A$ be a $\tau$-rigid $A$-module, then $^{\perp}\tau Z:=\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,\tau Z)=0\}$ is a torsion class of $\operatorname{\mathsf{mod}}\nolimits A$. The Bongartz completion $U_{Z}:=P(^{\perp}(\tau Z))$ of $Z$ is the maximal basic direct sum of $\operatorname{\mathsf{Ext}}$-projective objects in $^{\perp}(\tau Z)$, which is a $\tau$-tilting module. Let $B=\operatorname{\mathsf{End}}_{A}(U_{Z})$ and $C=B/\langle e_{Z}\rangle$, where $e_{Z}$ is the idempotent of $B$ corresponding to the projective $B$-module $\operatorname{\mathsf{Hom}}_{A}(U_{Z},Z)$. The algebra $C$ is called the $\tau$-tilting reduction of $A$ with respect to $Z$, see [J].
Before stating the main result of this subsection. Let us recall a result from [J]. Let $\mathcal{T}$ be a $k$-linear $\operatorname{\mathsf{Hom}}$-finite Krull-Schmidt triangulate category with a silting object $T$. Let $A=\operatorname{\mathsf{End}}_{\mathcal{T}}(T)$, and denote
@@
Then $T_{D}=X_{D}\oplus D$ lies in 2T-$\operatorname{\mathsf{silt}}\mathcal{T}:=\{T\in\operatorname{\mathsf{silt}}\mathcal{T}~|~T\in\operatorname{\mathsf{add}}\nolimits T*\operatorname{\mathsf{add}}\nolimits(T[1])\}$ (see [J, Proposition 4.9]). Jasso [J] has established the following connection between silting reduction and $\tau$-tilting reduction.
@@
Now we are in the position to state the main result of this subsection.
@@
Let $A$ be an endomorphism algebra of a basic silting complex over a hereditary abelian category and $Z$ be a $\tau$-rigid $A$-module.
@@
    The $\tau$-tilting reduction of $A$ with respect to $Z$ is an endomorphism algebra of a basic silting complex over a hereditary abelian category.
@@
    If $A$ is a quasi-silted algebra, then the $\tau$-tilting reduction of $A$ with respect to $Z$ is also a quasi-silted algebra.
@@
[…]
For the statement $(b)$. Let assume moreover $T$ is a 2-term silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$. Consider the following triangle in ${\mathcal{D}}^{b}(\mathcal{H})$
@@
It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(S_{T},M[i])=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{T},M[i])=0$ for $i\neq 0,1$. Hence $S_{N}$ is a 2-term silting complex over $D^{\perp_{\mathbb{Z}}}$. This completes the proof. ∎
@@
Let $A$ be a finite dimensional algebra over an algebraically closed field, and $Z$ be a $\tau$-rigid $A$-module. By Theorem 3.1 and Theorem 3.9 (b), if $A$ is shod, then the $\tau$-tilting reduction $\operatorname{\mathsf{End}}_{A}(U_{Z})/\langle e_{Z}\rangle$ is also a shod algebra.
@@ 4. Idempotent quotients for several classes of algebras
Let $A$ be a finite-dimensional $k$-algebra and $e$ an idempotent element in $A$.
@@
For an artin algebra $A$, denote by $\operatorname{\mathsf{ind}}A$ the set of isomorphism classes of indecomposable right $A$-modules. Given indecomposable $A$-modules $M$ and $N$, a path from $M$ to $N$ is a sequence
@@
where every $X_{i}\in\operatorname{\mathsf{ind}}A$ and every $f_{i}$ is a nonzero non-isomorphic morphism. The number $s$ is called the length of this path. Following [HRS], we have two full subcategories
@@
Let $e$ be an idempotent in $A$, for a right $A$-module $M$, we have a short exact sequence $0\to MeA\to M\to M/MeA\to 0$, where $MeA\in\operatorname{\mathsf{Fac}}eA$ and $M/MeA\in(eA)^{\perp}$.
@@
Consider a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$, $f:Q\to M$. Then we have a short exact sequence $0\to\operatorname{\mathsf{im}}\nolimits f\to M\xrightarrow{g}M/\operatorname{\mathsf{im}}\nolimits f\to 0$. Then $\operatorname{\mathsf{im}}\nolimits f\in\operatorname{\mathsf{Fac}}eA$. Set a morphism $h:eA\to M/\operatorname{\mathsf{im}}\nolimits f$. Since $eA$ is projective, $h$ is factorized by $g$, say $h=gu$ for some $u:eA\to M$. Then $u$ is factorized by $f$, since $f$ is a right $\operatorname{\mathsf{add}}\nolimits eA$-approximation. Then $u=fv$ for some $v:eA\to Q$. Hence $h=gu=gfv=0$. This means $M/\operatorname{\mathsf{im}}\nolimits f\in(eA)^{\perp}$.
Next we show $\operatorname{\mathsf{im}}\nolimits f=MeA$. Note that $\operatorname{\mathsf{im}}\nolimits f$ is independent of the choice of the right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation $f$. Hence firstly we construct a certain right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$. Assume $\{x_{1}e,\cdots,x_{d}e\}$ is a basis of the $k$-vector space $Me$. For each $1\leq i\leq d$, denote by $(x_{i}e)\cdot$ the left multiplication map $eA\to M$, $ea\mapsto x_{i}ea$, which is a right $A$-morphism. Then we have a morphism $\phi:(eA)^{d}\to M$ by the universal property of direct sum. By the isomorphism $Me\cong\operatorname{\mathsf{Hom}}_{A}(eA,M)$, $\{(x_{1}e)\cdot,\cdots,(x_{d}e)\cdot\}$ is a basis of $\operatorname{\mathsf{Hom}}_{A}(eA,M)$. Hence $\phi$ is a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$. In particular, $\operatorname{\mathsf{im}}\nolimits f=\operatorname{\mathsf{im}}\nolimits\phi$. By the definition of $\phi$, for any $\{ea_{i}\}_{i}\in(eA)^{d}$, $\phi(\{ea_{i}\}_{i})=x_{1}ea_{1}+\cdots+x_{d}ea_{d}$. This means $\operatorname{\mathsf{im}}\nolimits\phi\subseteq MeA$. On the other hand, $MeA$ is generated by elements in form of xeaxea with $x\in M$, $a\in A$. Set $xe=\sum\limits_{i=1}^{d}\lambda_{i}x_{i}e$, $\lambda_{i}\in k$. Then $xea=\sum\limits_{i=1}^{d}x_{i}e(\lambda_{i}a)\in\operatorname{\mathsf{im}}\nolimits\phi$. Hence $MeA=\operatorname{\mathsf{im}}\nolimits\phi=\operatorname{\mathsf{im}}\nolimits f$. This finishes the proof. ∎
@@
The projective $A$-module eAeA induces a torsion pair $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$ of $\operatorname{\mathsf{mod}}\nolimits A$. The sequence in Lemma 4.4 is the canonical sequence for $M$ with respect to $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$.
@@
For $M\in(eA)^{\perp}$, M is $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$, if and only if there is a projective $A$-module $P$ such that $M\cong P/PeA$.
@@
For the standard duality $\operatorname{\mathsf{D}}\nolimits=\operatorname{\mathsf{Hom}}_{k}(-,k)$, we have the following properties:
@@
    For an $A$-module $X$, we have $\operatorname{\mathsf{id}}_{A}X=\operatorname{\mathsf{pd}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(X)$ and $\operatorname{\mathsf{pd}}_{A}X=\operatorname{\mathsf{id}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(X)$. In particular, $A$ is shod if and only if $A^{op}$ is shod.
@@
    $(A/\langle e\rangle)^{op}=A^{op}/\langle e\rangle$.
@@
    For a $A$-module $X$, $X\in(eA)^{\perp}$ if and only if $\operatorname{\mathsf{D}}\nolimits(X)\in(Ae)^{\perp}$. In particular, the restriction of $\operatorname{\mathsf{D}}\nolimits$ to $(eA)^{\perp}$ is a duality, $\operatorname{\mathsf{D}}\nolimits:(eA)^{\perp}\to(Ae)^{\perp}$.
@@
    For a $A$-module $X\in(eA)^{\perp}$, it is $\operatorname{\mathsf{Ext}}$-projective (resp. $\operatorname{\mathsf{Ext}}$-injective) in $(eA)^{\perp}$ if and only if $\operatorname{\mathsf{D}}\nolimits(X)$ is $\operatorname{\mathsf{Ext}}$-injective (resp. $\operatorname{\mathsf{Ext}}$-projective) in $(Ae)^{\perp}$.

### v3

@@ 3.1. Idempotent quotients
For a triangulated category $\mathcal{T}$ and a subcategory $\mathcal{X}$, we denote by $\mathcal{X}]$ the ideal of $\mathcal{T}$ consisting of morphisms that factor through objects in $\mathcal{X}$.
@@
Note that $\langle e_{D}\rangle$ consists of those morphisms in $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ that factor through objects in $\operatorname{\mathsf{add}}\nolimits D$. In other words, $\langle e_{D}\rangle=[\operatorname{\mathsf{add}}\nolimits D](T,T)$. Let $f\in[\operatorname{\mathsf{thick}}D](T,T)$, and assume $f$ factors through $X\in\operatorname{\mathsf{thick}}D$ as $T\xrightarrow{a}X\xrightarrow{b}T$. By Lemma 2.4, we have
@@
Let $A$ be the endomorphism algebra of a basic silting complex over a hereditary abelian category, and $e\in A$ be an idempotent.
@@
    Then $A/\langle e\rangle$ is isomorphic to the endomorphism algebra of a silting complex over some hereditary abelian category.
@@
    Moreover, if $A$ is quasi-silted, then $A/\langle e\rangle$ is also quasi-silted.
@@
Suppose $\mathcal{H}$ is a hereditary abelian category and $T\in{\mathcal{D}}^{b}(\mathcal{H})$ is a basic silting object with $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$. Let $T=D\oplus N$ and $e_{D}$ be the idempotent of $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$ corresponding to $D$. It suffices to show that $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)/\langle e_{D}\rangle$ is isomorphic to the endomorphism algebra of a basic silting complex over some hereditary abelian category.
@@
where $f$ is a minimal right $\mathcal{S}$-approximation of $N$. It follows that $S_{N}\in\mathcal{S}^{\perp}$ by Lemma 2.2. On the other hand, by Lemma 2.4, one observes that $S\in\mathcal{S}_{D}^{\leq 0}:=\bigcup_{l\geq 0}\operatorname{\mathsf{add}}\nolimits D\ast\cdots\ast\operatorname{\mathsf{add}}\nolimits D[l]$.
@@
Now assume $T$ is 2-term (i.e. , $T$ satisfies (S1)) in ${\mathcal{D}}^{b}(\mathcal{H})$. In view of Proposition 2.8, we identify ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ with $\mathcal{S}^{\perp}$ via the embedding $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\hookrightarrow{\mathcal{D}}^{b}(\mathcal{H})$. Consequently, it remains to verify that $S_{N}$ is 2-term in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ with respect to $D^{\perp_{\mathbb{Z}}}$.
@@
Let $A$ be a finite-dimensional algebra over an algebraically closed field, and $e\in A$ be an idempotent. By Theorem 3.1 and Theorem 3.3 (b), if $A$ is shod, then $A/\langle e\rangle$ is also shod.
@@ 3.2. Idempotent subalgebras
Let $A$ be the endomorphism algebra of a basic silting complex over a hereditary category, and $e\in A$ be an idempotent.
@@
    The subalgebra $eAe$ is isomorphic to the endomorphism algebra of a basic silting complex over some hereditary category.
@@
    If $A$ is quasi-silted, then $eAe$ is also quasi-silted.
@@
    If $A$ is quasi-tilted, then $eAe$ is also quasi-tilted.
@@
(a) Let $A=\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$, where $\mathcal{H}$ is a hereditary abelian category, and $T$ is a basic silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$. Assume $T=U\oplus N$, and let $e=e_{N}$ be the idempotent corresponding to $N$. We first observe that $e_{N}(\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T))e_{N}\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N)$. By Theorem 2.17, there exists a presilting complex $D$, such that $D\oplus N$ is a basic silting object and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$. In addition, $N$ is a silting object in $(\operatorname{\mathsf{thick}}D)^{\perp}$. According to Proposition 2.7, $D^{\perp_{\mathbb{Z}}}$ is a hereditary abelian category. Furthermore, by Proposition 2.8, we identify $(\operatorname{\mathsf{thick}}D)^{\perp}$ with ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. Then $N$ is a basic silting complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$, and $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N)\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{D^{\perp_{\mathbb{Z}}}})}(N)$.
@@
Let $A$ be a finite-dimensional algebra over an algebraically closed field, and $e\in A$ be an idempotent. By Theorem 3.1 and Theorem 3.5 (b), if $A$ is shod, then $eAe$ is also shod.
@@
The statements in parts (b) and (c) of Theorem 3.5 were established in [AC]. Our method is different from the one in [AC].
@@
Let $Z\in\operatorname{\mathsf{mod}}\nolimits A$ be a $\tau$-rigid $A$-module, then $^{\perp}(\tau Z):=\{X\in\operatorname{\mathsf{mod}}\nolimits A\mid\operatorname{\mathsf{Hom}}_{A}(X,\tau Z)=0\}$ is a torsion class of $\operatorname{\mathsf{mod}}\nolimits A$. The Bongartz completion $U_{Z}:=P(^{\perp}(\tau Z))$ of $Z$ is the maximal basic direct sum of $\operatorname{\mathsf{Ext}}$-projective objects in $^{\perp}(\tau Z)$, which is a $\tau$-tilting module. Let $B=\operatorname{\mathsf{End}}_{A}(U_{Z})$ and $C=B/\langle e_{Z}\rangle$, where $e_{Z}$ is the idempotent of $B$ corresponding to the projective $B$-module $\operatorname{\mathsf{Hom}}_{A}(U_{Z},Z)$. The algebra $C$ is called the $\tau$-tilting reduction of $A$ with respect to $Z$, see [J].
Before stating the main result of this subsection. Let us recall a result from [J]. Let $\mathcal{T}$ be a $k$-linear $\operatorname{\mathsf{Hom}}$-finite Krull-Schmidt triangulated category with a silting object $T$. Let $A=\operatorname{\mathsf{End}}_{\mathcal{T}}(T)$, and denote
@@
Then $T_{D}:=X_{D}\oplus D$ lies in 2T-$\operatorname{\mathsf{silt}}\mathcal{T}:=\{T\in\operatorname{\mathsf{silt}}\mathcal{T}~|~T\in\operatorname{\mathsf{add}}\nolimits T*\operatorname{\mathsf{add}}\nolimits(T[1])\}$ (see [J, Proposition 4.9]). Jasso [J] has established the following connection between silting reduction and $\tau$-tilting reduction.
@@
Now we are in a position to state the main result of this subsection.
@@
Let $A$ be the endomorphism algebra of a basic silting complex over a hereditary abelian category, and $Z$ be a $\tau$-rigid $A$-module.
@@
    The $\tau$-tilting reduction of $A$ with respect to $Z$ is isomorphic to the endomorphism algebra of a basic silting complex over some hereditary abelian category.
@@
    If $A$ is quasi-silted, then the $\tau$-tilting reduction of $A$ with respect to $Z$ remains a quasi-silted algebra.
@@
Let $\mathcal{H}$ be a hereditary abelian category and $T\in{\mathcal{D}}^{b}(\mathcal{H})$ a basic silting object with $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}{\mathcal{H}}}(T)$. According to [IJY, Theorem 4.5], there is a bijection
$2_{T}\text{-}\operatorname{\mathsf{silt}}{\mathcal{D}}^{b}(\mathcal{H})\leftrightarrow s\tau\text{-}\operatorname{\mathsf{tilt}}A$
induced by $\overline{(-)}=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,-)$. Thus, there is a presilting complex $D$ in ${\mathcal{D}}^{b}(\mathcal{H})$ contained in $\operatorname{\mathsf{add}}\nolimits T*\operatorname{\mathsf{add}}\nolimits T[1]$ such that $Z=\overline{D}$. By Theorem 3.8, we have $\operatorname{\mathsf{End}}_{A}(U_{Z})/\langle e_{Z}\rangle\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}}(T_{D})$, where $\mathcal{S}=\operatorname{\mathsf{thick}}D$. According to Proposition 2.13, we have ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\cong{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}$. On the other hand, by the definition of $T_{D}$, $T_{D}\cong T$ in ${\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}$. It follows that $T_{D}$ is a basic silting object of ${\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}$ by Theorem 2.5. Consequently, $\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})/\mathcal{S}}(T_{D})$ is isomorphic to the endomorphism algebra of a basic silting complex over $D^{\perp_{\mathbb{Z}}}$. This finishes the proof of $(a)$.
To establish part (b), we further assume that $T$ is a 2-term silting complex in ${\mathcal{D}}^{b}(\mathcal{H})$. Consider the following triangle in ${\mathcal{D}}^{b}(\mathcal{H})$
@@
It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})}(S_{T},M[i])=\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(S_{T},M[i])=0$ for $i\neq 0,1$. Hence $S_{N}$ is a 2-term silting complex over $D^{\perp_{\mathbb{Z}}}$. This completes the proof.
@@
Let $A$ be a finite-dimensional algebra over an algebraically closed field, and $Z$ be a $\tau$-rigid $A$-module. By Theorem 3.1 and Theorem 3.9 (b), if $A$ is shod, then the $\tau$-tilting reduction $\operatorname{\mathsf{End}}_{A}(U_{Z})/\langle e_{Z}\rangle$ is also a shod algebra.
@@ 4. Idempotent quotients for several classes of algebras
Let $A$ be a finite-dimensional $k$-algebra and $e\in A$ an idempotent.
@@
Remark 4.3.
It is worth noting that while the classes of shod, laura, and weakly shod algebras are closed under taking arbitrary idempotent quotients, the same does not hold for tilted or quasi-tilted algebras. As evidenced by the example in [BMR, Section 2.5], a tilted algebra can yield a quotient that is strictly shod (i.e., shod but not quasi-tilted).
Let $A$ be an Artin algebra. Given indecomposable $A$-modules $M$ and $N$, a path from $M$ to $N$ is a sequence of morphisms
@@
where each $X_{i}\in\operatorname{\mathsf{ind}}A$ and each $f_{i}$ is a nonzero non-isomorphism. The integer $s$ is the length of this path. Following Happel, Reiten, and Smalø [HRS], we define two full subcategories of $\operatorname{\mathsf{ind}}A$
@@
Let $e\in A$ be an idempotent. For any right $A$-module $M$, there exists a short exact sequence $0\to MeA\to M\to M/MeA\to 0$, where $MeA\in\operatorname{\mathsf{Fac}}eA$ and $M/MeA\in(eA)^{\perp}$.
@@
Consider a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$, denoted by $f:Q\to M$. This induces the short exact sequence:
$0\to\operatorname{\mathsf{im}}\nolimits f\to M\xrightarrow{g}M/\operatorname{\mathsf{im}}\nolimits f\to 0,$
where it is immediate that $\operatorname{\mathsf{im}}\nolimits f\in\operatorname{\mathsf{Fac}}eA$. Let $h:eA\to M/\operatorname{\mathsf{im}}\nolimits f$ be any morphism. Since $eA$ is projective, $h$ factors through $g$, say $h=gu$ for some $u:eA\to M$. Furthermore, as $f$ is a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation, the morphism $u$ factors through $f$, implying $u=fv$ for some $v:eA\to Q$. Hence $h=gu=gfv=0$. This means $M/\operatorname{\mathsf{im}}\nolimits f\in(eA)^{\perp}$.
We now verify that $\operatorname{\mathsf{im}}\nolimits f=MeA$. First, observe that the image of a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation is independent of the specific choice of $f$. To construct an explicit approximation, let $\{x_{1}e,\dots,x_{d}e\}$ be a $k$-basis for the vector space $Me$. For each $1\leq i\leq d$, let $\theta_{i}:eA\to M$ denote the right $A$-morphism defined by left multiplication: $ea\mapsto x_{i}ea$. By the universal property of direct sums, these maps induce a morphism $\phi:(eA)^{d}\to M$. Under the canonical isomorphism $Me\cong\operatorname{Hom}_{A}(eA,M)$, the set $\{\theta_{1},\dots,\theta_{d}\}$ forms a $k$-basis for $\operatorname{Hom}_{A}(eA,M)$. Consequently, $\phi$ constitutes a right $(\operatorname{\mathsf{add}}\nolimits eA)$-approximation of $M$, and it follows that $\operatorname{\mathsf{im}}\nolimits f=\operatorname{\mathsf{im}}\nolimits\phi$. To show the equality of submodules, note that for any element $(ea_{1},\dots,ea_{d})\in(eA)^{d}$, we have:
$\phi(ea_{1},\dots,ea_{d})=\sum_{i=1}^{d}x_{i}ea_{i}\in MeA,$
which implies $\operatorname{\mathsf{im}}\nolimits\phi\subseteq MeA$. Conversely, $MeA$ is generated by elements of the form xeaxea where $x\in M$ and $a\in A$. Expanding $xe$ in terms of our chosen basis, we write $xe=\sum_{i=1}^{d}\lambda_{i}x_{i}e$ for some $\lambda_{i}\in k$. Then:
$xea=\sum_{i=1}^{d}x_{i}e(\lambda_{i}a)\in\operatorname{\mathsf{im}}\nolimits\phi.$
Therefore, $MeA=\operatorname{\mathsf{im}}\nolimits\phi=\operatorname{\mathsf{im}}\nolimits f$, completing the proof. ∎
@@
The projective $A$-module eAeA induces a torsion pair $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$ of $\operatorname{\mathsf{mod}}\nolimits A$. The sequence in Lemma 4.5 is the canonical sequence of $M$ with respect to $(\operatorname{\mathsf{Fac}}eA,(eA)^{\perp})$.
@@
Let $M\in(eA)^{\perp}$. Then M is $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$ if and only if there is a projective $A$-module $P$ such that $M\cong P/PeA$.
@@
Let $\operatorname{\mathsf{D}}\nolimits=\operatorname{\mathsf{Hom}}_{k}(-,k)$ be the standard duality. The following properties hold:
@@
    For any $A$-module $X$, we have $\operatorname{\mathsf{id}}_{A}X=\operatorname{\mathsf{pd}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(X)$ and $\operatorname{\mathsf{pd}}_{A}X=\operatorname{\mathsf{id}}_{A^{op}}\operatorname{\mathsf{D}}\nolimits(X)$. Consequently, $A$ is shod if and only if its opposite algebra $A^{op}$ is shod.
@@
    $(A/\langle e\rangle)^{op}\cong A^{op}/\langle e\rangle$.
@@
    For an $A$-module $X$, $X\in(eA)^{\perp}$ if and only if $\operatorname{\mathsf{D}}\nolimits(X)\in(Ae)^{\perp}$. In particular, the restriction of $\operatorname{\mathsf{D}}\nolimits$ induces a duality $\operatorname{\mathsf{D}}\nolimits:(eA)^{\perp}\to(Ae)^{\perp}$.
@@
    For an $A$-module $\in(eA)^{\perp}, XX$ is $\operatorname{\mathsf{Ext}}$-projective (resp. $\operatorname{\mathsf{Ext}}$-injective) in $(eA)^{\perp}$ if and only if $\operatorname{\mathsf{D}}\nolimits(X)$ is $\operatorname{\mathsf{Ext}}$-injective (resp. $\operatorname{\mathsf{Ext}}$-projective) in $(Ae)^{\perp}$.